OpenStudy (anonymous):

Naomi has three pounds of apples and 2 1/2 pounds of grapes if she gets 1/3 of her apples to Christine, how many pounds of apples does she have left?

OpenStudy (capthca):

this is history plz go to math for that plz

OpenStudy (saucvi):

This is history but the answer is 1 1/6. You find common denominators which in this case would be 6. so 1/3 = 2/6 and 2 1/2 = 15/6. Now subtract 2/6 from 15/6 and you get 13/6. Simplified, 2 1/6.